Circular No. 6724

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997dd IN NGC 6060 S. Nakano, Sumoto, Japan, reports the discovery of an apparent super-

nova (V = 16.4) by Masakatsu Aoki (Tsukioka Cho, Toyama) on ten CCD frames (limiting mag ∼ 17.5) taken with a 0.43-m f/6 reflector around Aug. 20.535 UT. SN 1997dd is located at α = 16^{h}05^{m}46^{s}.32, δ = +21^{o}29′14′′.3 (equinox 2000.0), which is 79′′ west and 9′′ north of the center of NGC 6060. Confirming frames were taken by Aoki on Aug. 21, and nothing was visible at this location on his frames taken on May 26, June 14 and 27, July 21, and Aug. 3 and 10. P. Garnavich and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that K. Stanek obtained CCD images on Aug. 26.18 with the Whipple Observatory’s 1.2-m reflector, showing SN 1997dd at position end figures 46^{s}.33, 14′′.2 (± 0′′.4), and magnitudes R ∼ 17.2, B ∼ 17.9 (measured with respect to the USNO A1.0 magnitudes of the star at position end figures 52^{s}.60, 31′26′′.8).

CEPHEUS X-4 J. M. Bonnet-Bidaud, Service d’Astrophysique, Centre d’Etudes de

Saclay; and M. Mouchet, Observatoire de Paris, report: “Long-slit spectra (375–725 nm; resolution 1.3 nm) of the proposed optical identification of the transient 66-s x-ray pulsar Cep X-4 = GS 2138+56 (IAUC 6698, 6711), obtained on July 29 and 30 at Haute-Provence Observatory, show features typical of x-ray Be binaries. Superimposed on a weak emission from the IC 1396 nebula, strong Hα (equivalent width 4.5 nm) and Hβ (EW 0.3 nm) lines are seen in emission with the other Balmer lines in absorption. Significant interstellar absorption features are also detected, including a strong Na I doublet (589 nm) and diffuse bands at 443, 578, and 628 nm. Considering a wide range of spectral types from B7 to O9, dereddening of the spectra yields E_{B−V} = 1.0–1.2, corresponding to an absorption of N_{H} = (7–8) × 10^{21} atoms cm^{−2} for a typical gas/dust ratio. This is in very good agreement with the x-ray column density (7.8 ± 1.9) × 10^{21} atoms cm^{−2} derived by Schulz et al. (1995, A.Ap. 295, 413) and confirms the identification. Assuming a mean interstellar density of 1 hydrogen atom cm^{−3}, the deduced distance is 2300–2700 pc. A similar value of 3000 pc is independently derived from the strong Na line (EW 0.3 nm), if assumed fully interstellar. At this distance, the x-ray quiescent luminosity is 2 × 10^{33} erg s^{−1}, comparable to x-ray Be-binary low states.”

1997 August 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green